OpenStudy (the_legend_of_kris):

You're looking at a catalog that sells cooking supplies. Some of the items are drawn in scale at a ratio of 1:12, in cm. If a scale drawing of one of these items is 3 cm wide, how wide is the actual items in cm?

OpenStudy (lapinblanc):

Ratio solutions follow: a/b=c/d you have 1/12=3/x solve for x

OpenStudy (the_legend_of_kris):

How do I solve for X?

OpenStudy (the_legend_of_kris):

I believe its 36cm, but I'm not sure.

OpenStudy (meredith17):

Step 1: Cross-multiply. 1 12 = 3 x 1*x=(3)*(12) x=36 Answer: x=36

OpenStudy (lapinblanc):

That answer is exactly correct. To solve for x you would have multiplied by reciprocals, which would have looked like: 1/12=3/x 12(1/12)=12(3/x) 1=36/x x*1=x(36/x) x=36